OpenStudy (anonymous):

1- -40/5=9? is it correct? @perl

OpenStudy (anonymous):

@mathstudent55

OpenStudy (solomonzelman):

yes, you are correct.

OpenStudy (solomonzelman):

1 - (-40/5) =9 1 + (40/5)=9 1+8=9 9=9

OpenStudy (anonymous):

are you sure? can you explain it to me/

OpenStudy (solomonzelman):

I did above

OpenStudy (anonymous):

Thank you! @SolomonZelman

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

good work